Circular No. 7329

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999gi IN NGC 3184 S. Nakano, Sumoto, Japan, reports the discovery by Reiki Kushida,

Yatsugatake South Base Observatory, of a supernova (mag 14.5) on unfiltered CCD frames taken on Dec. 9.82 UT with a 0.40-m reflector. Previous frames that she took between 1994 and 1999 May 6 show nothing at this location (limiting mag 18), and nothing appears at the location of the new object on a Digital Sky Survey image. Position measured by Y. Kushida for SN 1999gi: α = 10^{h}18^{m}16^{s}.66, δ = +41^{o}26′28′′.2 (equinox 2000.0), which is 3′′.5 west and 60′′.5 north of the center of NGC 3184. W. Li, University of California at Berkeley, reports that a KAIT image of NGC 3184 taken on Dec. 2.5 shows nothing at the position of SN 1999gi (limiting mag 19.0).

S. Jha, P. Garnavich, P. Challis, and R. Kirshner, Harvard-Smithson-ian Center for Astrophysics, report that a spectrum of SN 1999gi, taken by A. Mahdavi and Jha with the F. L. Whipple Observatory (FLWO) 1.5-m telescope (+ FAST spectrograph) on Dec. 10.42 UT, reveals it to be a type-II supernova at an early epoch. The spectrum exhibits a blue continuum with a broad Hα P-Cyg profile, as well as a hint of He I (rest 587.6 nm) absorption. The photospheric expansion velocity, measured from narrow superimposed Hα emission to the trough of the Hα profile, is 13 300 km/s. The spectrum also exhibits a P-Cyg profile with absorption trough observed at 444.3 nm, and a slightly less broad absorption feature observed at 394.0 nm. In addition, interstellar Na I D with equivalent width = 0.096 nm is visible at the redshift of NGC 3184 (given by the NASA/IPAC Extragalactic Database as 532 km/s), implying that the supernova is likely significantly extinguished by dust in the host galaxy. Concurrent images of SN 1999gi, taken by J. Munoz Lozano at the FLWO 1.2-m telescope (+ 4Shooter), yield position end figures for the supernova of 16^{s}.62, 28′′.6.

V335 VULPECULAE U. Munari, T. Tomov, and D. Moro, Padua and Asiago Observatories;

and A. Henden, U.S. Naval Observatory, Flagstaff, write: “The carbon symbiotic star V335 Vul seems to be undergoing a bright outburst. Photometry with the USNO 1.0-m in Flagstaff: June 23.32 UT, V = 12.87, B − V = +5.01; Dec. 3.05, V = 11.31, B −V = +3.06. Echelle and Boller & Chivens spectroscopy obtained with the Asiago 1.22- and 1.82-m telescopes indicate a tenfold increase in the intensity of the emission lines, compared to quiescent conditions (cf. IBVS 4668). The strongest emission lines include Hα, Hβ, Na I D, and the near-infrared Ca II triplet.”

1999 December 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green